# An engineering guide to superconducting quantum circuit shielding

### v1

@@ II.2.3 Practical implementation of electromagnetic shielding
The magnetic field was simulated using the finite element method. Two cylindrical shields were considered, nested one inside the other. The shields were placed in a vacuum parallelepiped. The external magnetic field of 50 $\mu T$ along the axis of the cylinders was set. Relative magnetic permeabilities were assigned for the $\mu$-metal (70,000), superconductor ($1\cdot 10^{-6}$), and vacuum (1). Magnetic induction distribution in the area was simulated. The thickness of the shields is set to 1 $mm$. Outer diameters of the shields are 66 $mm$ and 70 $mm$. The heights are 180 $mm$ and 190 $mm$. The calculation scheme and simulation results are presented in Figure 14. In the double-shield configuration, the minimum wall-to-wall separation between shields was maintained at 2 $mm$.
Based on the modeling, the combination of aluminum shield inside and $\mu$-metal shield outside can provide a minimum level of magnetic field at the SQC position inside the shields. The combination of magnetic shields should be chosen based on the required level of the external magnetic field.
@@ II.4 Shielding configurations known at present
Figure 15: A review of shielding configurations. “SC” – superconducting layer, “IR” – infrared radiation absorbing layer, “${\mu}$” – high magnetic permeability layer, and “Refl.” – reflective infrared radiation layer. “Quantity” is the number of referenced papers employing each specific shielding configuration.
Figure 16: Shielding systems with various IR shielding placement: a) a 2-layer shielding: copper shield with an IR-absorber layer and cryoperm shield Eddins et al. (2018); b) a 4-layer shielding with two IR shields and two magnetic shields (cryoperm and superconducting) Burnett et al. (2019); c) a 3-layer shielding with aluminum foil IR shield and mu-metal shield Schwartz (2016).
Let’s consider and analyze the existing shielding systems according to the previously discussed recommendations. Various configurations of multilayered shielding systems are illustrated in Fig. 15. Two cases are presented in histogram: in the first case the sample holder is considered as a part of the shielding, and in the second it is not. The shielding configurations in Fig. 15 represent the order of shields used: starting from the sample holder (or the first shield) and finishing with the last shield within the system.
@@
We would like to emphasize that the frequency of shielding systems mention in literature does not necessarily correlate with optimal performance. The subsequent analysis in this section aims to justify the prevalence of specific shielding configurations in the literature based on the presented in this paper simulation results.
Shielding from IR is predominantly accomplished with a single layer. Two layers of absorbent coatings are used in only 10% of cases Kroll (2019); Liu (2016a); Bengtsson (2020), provided that we do not take into account the sample holder. The rationality of using just one shielding layer is also supported by our modeling results (see Fig. 6). However, taking into account absorbent coatings on the sample holder, we can find that most shielding systems employ two IR layers (Fig. 15). According to our calculations, the sample holder inner surface plays a significant role in protecting SQC from IR radiation. To simplify the shielding system design, an absorbing coating can be applied only on the inner surface of the sample holder, without using an additional IR shield around.
Figure 17: Shielding systems with various magnetic and superconducting shielding configurations: a) a 2-layer shielding without an IR-absorber Yan et al. (2016); b) a double-layered cryoperm shielding Schmitt (2015); c) a 4-layer shielding: a superconducting sample holder, lead and two $\mu$-metal shields Burnett (2013).
@@
Approximately in 80% of considered works IR shielding is located around the sample holder (Fig. 16a) Schwartz (2016); Pop et al. (2014); Gyenis et al. (2021); Jin et al. (2015); Serniak (2019). In other cases it surround the coldest plate of the fridge as shown in Fig. 16b Burnett et al. (2019); Bengtsson (2020). IR shields are typically positioned immediately after the sample holder Poletto et al. (2012); Serniak (2019); Roch et al. (2014); Dickel (2018). However, in multilayered shielding systems, they can be located at second (Fig. 16c) Burnett et al. (2019) or even fourth Grünhaupt et al. (2017); Bengtsson (2020) shielding layer. Modeling results show that the multilayer IR shielding have the same effect as a single-layer one.
Only 12% of superconducting shields have two layers Bengtsson (2020); Anton et al. (2013); Sendelbach (2013). This is well explained by the simulation results, according to which the number of superconducting shielding layers does not affect the shielding quality. The authors in Inc. state that the use of a superconducting shield with a ${\mu}$-metal shield can achieve a magnetic field level of approximately 1 $nT$. It is also supported by the modeling results presented above. The provided magnetic field is sufficient to protect the SQC from external magnetic field fluctuations. From Fig. 15, it can be seen that the majority of superconducting shielding is realized by the sample holder, what is another way for simplify the shielding system.
@@
The location of the superconducting shields may vary, most of them directly surround the holder Jin et al. (2015); Murch et al. (2013); Shankar et al. (2013), while the others located either at lower Burnett et al. (2019); Bengtsson (2020) or penultimate stage of the fridge Kreikebaum et al. (2016). Due to the requirement of the magnetic field to be uniformly distributed at the sample holder place, it is preferable to close the superconducting shield by the lid.
@@
Superconducting shields may be located either right after the holder (Fig. 17a) Yan et al. (2016); Liu (2016b); Sendelbach (2013) or after IR shield Serniak (2019); Murch et al. (2013); Dickel (2018). In systems with more than four shields, a superconducting shield may be placed at third layer from the holder Liu (2016a), or even at fifth Burnett et al. (2019); Bengtsson (2020). However, our modeling results showed that multilayered shielding system do not significantly increase shielding efficiency, so a more compact design may be used.
In approximately in 80% of considered works, single-layer magnetic shielding was used, in 19% two-layer shielding was used Burnett (2013); Yeh et al. (2017); Asaad et al. (2016); Tabuchi et al. (2016), and only in one work Buks et al. (2020) presented three-layer shielding was used. As it was previously demonstrated, the combination of a superconducting shield inside and a ${\mu}$-metal shield outside provides the most effective protection against magnetic fields.
@@
As well as the other types of the shields, magnetic shields are typically placed around the holder Burnett et al. (2019); Rigetti et al. (2012); Kelly (2015); Mirhosseini et al. (2019); Sung et al. (2021). At the same time, there are more options for the location of magnetic shields within the fridge, such as the lower stage (Fig. 17b) Jin et al. (2015); Mirhosseini et al. (2020), 2.7 $K$ Kreikebaum et al. (2016) or 4 $K$ Barends et al. (2011); Cooper (2013), or around the fridge itself (Fig. 17c) Córcoles et al. (2011); Moghaddam et al. (2019); Sendelbach (2013).
@@
How it was been determined in the modeling, there is requirement for the shield to be closed, so it may only be placed around the sample holder. In case the magnetic shield is the only one shield used in a system, it typically follows right after the sample holder Vidiborskiy et al. (2013); Li et al. (2021); Pierre et al. (2014). In other cases, the magnetic shield may be placed outside of the other shields, such as IR shields Chou (2018); Wu et al. (2021); Roch et al. (2014) and superconducting shields Monroe et al. (2021a); Asaad et al. (2016); Sendelbach (2013). However, in case of multilayered magnetic shielding system, ${\mu}$-metal shields may be inserted between the superconducting shields Burnett et al. (2019); Bengtsson (2020). As it was previously mentioned, for optimal shielding performance, the superconducting shield should be placed inside the ${\mu}$-metal one.
@@
Most considered shielding systems meet our requirements for the shielding. Requirements description for the shields based on the modeling results make it possible to identify excessiveness in the shielding system and make it lighter and more compact. We also mark on Fig. 15 (by a green frame) the best shielding systems based on our simulation results. If it is possible to use all three types of shields, we recommend to use “IR + SC + $\mu$-metal” system. If, due to some limitations, it is possible to use only two shields, we recommend the system “IR+SC”, the less preferred configuration is “IR+$\mu$-metal”.
@@ III MICROWAVE LINES FILTERING
For high performance of SQC it is necessary to filter the frequencies out of band for all the types of lines listed above. Insufficient filtering of frequencies close to the qubit ones can lead to parasitic single-qubit rotations, degrading the single qubit gate fidelities, as well as undesirable population of possible coupled with qubits parasitic modes, which can lead to additional qubits decoherence. Insufficient filtering of IR frequencies (above 90 $GHz$) can lead to cooper-pairs breaking in superconductor thin film and also degrade the qubit coherence. In this article we focus only on the different filter types used for all lines listed above. The purpose of other microwave components, used in control lines, such as attenuators, circulators, isolators etc., is described in detail in Ref.Krinner et al. (2019).
@@ III.1 Requirements for filters
Filtering is usually used to protect SQCs from stray electromagnetic fields and IR radiation that can propagate in microwave lines. For this purpose, various types of radio frequency (RF) and microwave (MW) filters (Low-pass filter - LPF Premkumar et al. (2021); Arute et al. (2019); Serniak (2019); Moghaddam et al. (2019); Bengtsson et al. (2020), High-pass filter - HPF Burnett (2013); Kudra et al. (2020); Wisby (2017) or Band-pass filter - BPF Buks et al. (2020); Lisenfeld et al. (2019); Mundhada (2020)) covering the SQC operating range (typically, 4-8 GHz) are usually used. RF and MW filters are designed to cut off a portion of the signal spectrum outside the required passband, suppressing most of the stray electromagnetic fields. For suppressing IR radiation, a special IR filters are usually used. The working principle of such filters based on IR radiation power loss due to eddy currents excitation in filter filler material Wang et al. (2014).
It is necessary to use RF and MW filters with IR ones, as they cut-off different parts of the signal spectrum. For this purpose, one can use BPF+IR or a combination of HPF+LPF+IR. However, in considered measurement circuits, it is more convenient to use either a single LPF or a single IR filter, or various combinations of these filters. The main requirement for IR filters is as low as possible attenuation in the SQC operating frequency range, smooth transmission characteristic and at the same time high enough attenuation (no less than 20 $dB$) at higher frequencies (above 20 $GHz$).
@@ III.2 IR filters
In works, IR filters may have different designations (IR filter Chou (2018); Opremcak (2020); Sank (2014); Schneider et al. (2019), lossy filter Schwartz (2016); Murch et al. (2013); Tan et al. (2015), thermalization filter Kreikebaum et al. (2016), etc.), but all of them are designed for the same purpose: to suppress stray IR radiation in signal lines. The filling material for such filters (see Fig. 18) typically consists of an epoxy resin containing fine particles of carbonyl iron (Eccosorb CR-110 Place et al. (2021); Narla (2017); Reagor (2015); Ivanov et al. (2023), or, less frequently, CR-124 Krinner et al. (2019)), and metal powder Barends et al. (2011) (mainly copper Megrant et al. (2012); Cooper (2013); Colless et al. (2018)). Other possible fillers include epoxy resin based on aluminum oxide (Stycast) mixed with bronze powder de Visser (2014), or without it Lisenfeld et al. (2019). Specific fillers such as carbon nanotubes may also be used Moghaddam et al. (2019).
There are no studies with a direct comparison of the IR filters efficiency with different types of filling material. However, in considered works, Eccosorb CR-110 epoxy is the most widely used in IR filters (see Fig. 18).
@@
IR filters are employed in all types of signal lines of cryogenic setup. In the following paragraphs we will analyze various filtering configurations.
III.3 Filtering of input readout line
Typically, in cryogenic setups one Córcoles et al. (2015); Wu et al. (2021); Buks et al. (2020); Martin et al. (2020); Lachance-Quirion et al. (2017) or two filters Gyenis et al. (2021); Shankar et al. (2013); Sank (2014); Schneider et al. (2019); Opremcak et al. (2021) are used at the input readout line. Less frequently, three filters are installed in a row Chen (2018); Colless et al. (2018); Grünhaupt et al. (2019). RF and MW and IR filters can be used both individually Schwartz (2016); Yan et al. (2016); Grünhaupt et al. (2017); Andersen et al. (2019) and together Zhang (2018); Lisenfeld et al. (2019); Leonard et al. (2019). Different combinations of filters for the input readout line are shown in Fig. 19.
In most measurement circuits, filters are located outside the shielding Jin et al. (2015); Arute et al. (2019); Yeh et al. (2017); Jr. (2018); Kudra et al. (2020); Wilen et al. (2021). Less frequently, they may be located both inside and outside the shielding Burnett et al. (2019); Premkumar et al. (2021); Gao (2018); Mundhada (2020). In very rare cases, the filters are only located inside the shielding Kreikebaum et al. (2016), mounted on the lid of the shield Hazard (2019); Barends et al. (2011), or just prior to the sample holder Cooper (2013). In Serniak (2019) it is indicated that the location of an IR filter inside the shielding can reduce the quasiparticles generation in a qubit. Thus, it is recommended to place the IR filter within a metal shield thermalized on the base stage of the dilution refrigerator. For RF and MW filters there is no special requirements for placement in the measurement circuit. Thus, they are typically located outside the shielding.
As it was mentioned before, it is better to use RF and MW filters and IR ones together, the number of filters at the input readout line should be at least two. The recommended configurations should be the following: either BPF + IR or HPF + LPF + IR. For the filler of IR filters, Eccosorb CR-110 epoxy is the most popular material used in considered works.
Figure 19: A review of filtering configurations for readout (RO), drive, and flux lines. LPF stands for a low pass filter, HPF for a high pass filter, and BPF for a band pass filter. IR represents an infrared radiation filter. The “Quantity” represents the number of papers (from references list), which use the exact filtering type.
III.4 Filtering of output readout line
The output readout line experiences the same types of radiation effects, as the input line. Therefore, for this line the same recommendations can be made as for the input readout line. For the output line, a little more variety of filter configurations are used. For instance, at the output readout line a combination of HPFs and BPFs are frequently used (see Fig. 19).
Typically, two filters are used, but there are also single filters or combinations of three different filters. As shown in Fig. 1, the filters are generally placed at the lower stage of the dilution refrigerator. In addition, we recommend to place only IR filters inside the shield thermalized on the coldest stage of the cryostat. The design and materials used for IR filters at the output line are the same as for the input one.
III.5 Drive line filtering (XY control)

### v2

@@ II.2.3 Practical implementation of electromagnetic shielding
The magnetic field was simulated using the finite element method. Two cylindrical shields were considered, nested one inside the other. The shields were placed in a vacuum parallelepiped. The external magnetic field of 50 $\mu T$ along the axis of the cylinders was set. Relative magnetic permeabilities were assigned for the $\mu$-metal (70,000), superconductor ($1\cdot 10^{-6}$), and vacuum (1). Magnetic induction distribution in the area was simulated. The thickness of the shields is set to 1 $mm$. The heights of the shields are 180 $mm$ and 190 $mm$. Outer diameters are 66 $mm$ and 70 $mm$. In the double-shield configuration, the minimum wall-to-wall separation between shields was maintained at 2 $mm$. The calculation scheme and simulation results are presented in Figure 12. The simulation of magnetic shield configurations yields the magnetic field distribution inside the shielding.
For quantitative comparison of different shielding configurations, we calculated magnetic field suppression (MFS) factor within the shielded volume, defined as:
$MFS=\mid 20lg\frac{B_{average1/3}}{B_{external}}\mid dB,$ (4)
where $B_{external}$ – external magnetic field 50 $\mu T$; $B_{average1/3}$ – the mean magnetic field value within the lower third section of the shields (where the sample holder is typically positioned).
Modeling results indicate that the optimal configuration for minimizing magnetic field at the SQC position consists of an inner aluminum shield and outer $\mu$-metal shield. When the aluminum shield is positioned inside the $\mu$-metal shield, it operates within a pre-attenuated magnetic field, yielding enhanced overall attenuation. This improvement arises from the complementary shielding mechanisms: $\mu$-metal "pulls over" magnetic fields through its high magnetic permeability, while superconducting materials exclude magnetic fields via the Meissner effect.
Placing the superconducting shield externally causes magnetic field lines to distort around the outer shield and concentrate near the opening, potentially compromising sample protection. In contrast, the $\mu$-metal outer configuration provides superior attenuation by absorbing a substantial portion of incident fields before the superconducting shield expels residual flux. An additional consideration is that external placement of superconducting shields may lead to flux trapping phenomena.
The combination of magnetic shields should be chosen based on the required external magnetic field suppression level. Regarding target field determination, the required magnetic field suppression should be based on the quantum circuit dimensions. For illustration, consider a frequency-tunable transmon qubit with a SQUID loop area ($S_{SQUID}$) of 30 $\mu m^{2}$. The magnetic flux quantum of $\Phi_{0}=2.07\cdot 10^{-15}Wb$ establishes a reference field:
$\Phi=B\cdot dS\Rightarrow B_{0}=\frac{\Phi_{0}}{S_{SQUID}}=\frac{2.07\cdot 10^{-15}}{30\cdot 10^{-12}}\approx 68\mu T$ (5)
This calculation indicates that approximately 69 $\mu$T external field would induce one flux quantum in such a SQUID. Consequently, effective shielding should reduce ambient fields to 0.1-1% of this value (0.07-0.69 $\mu$T) to ensure residual fields remain negligible and compensable via flux-bias lines. For this particular case, shielding with MFS of at least 60 $dB$ will be required, necessitating either a system with a minimum of two $\mu$-metal shields, or a single superconducting shield.
@@
Modeling results demonstrate that open cylindrical shielding systems without lids achieve minimal magnetic field strength in their lower third section. When lid implementation is problematic, optimal shielding performance requires positioning the sample holder as close as possible to the bottom of the shielding system.
@@ II.4 Shielding configurations known at present
Figure 13: A review of existing shielding configurations: a) magnetic field suppression efficiency for shielding systems on the diagram below; b) symbols on the diagram: “Cu” – copper, “?” – not indicated, “/” – separation between the sample holder and shielding system,“SC” – superconducting layer, “IR” – infrared radiation absorbing layer, “$\mu$” – high magnetic permeability layer, and “Refl.” – reflective infrared radiation layer. “Quantity of papers” is the number of referenced papers employing each specific shielding configuration. The color mapping of shielding configurations represents both magnetic field suppression efficiency and system complexity: green indicates configurations achieving high field suppression with relatively simple designs, while red denotes systems that provide comparable suppression but require substantially more complex implementations. Additionally, the column coloring scheme identifies the positioning of IR-absorbing layers within each shielding system.
Let’s consider and analyze the existing shielding systems according to the previously discussed recommendations. Various configurations of multilayered shielding systems are shown in Figure 13. Figure 13 illustrates the hierarchical organization of shielding configurations, presented in sequential order from the sample holder (before the slash) to the innermost shield (immediately after the slash), culminating with the outermost shield in the system. The upper portion of the diagram (Figure 13a) displays the magnetic field suppression efficiency for various shielding systems shown in the lower portion (Figure 13b). These efficiency metrics were derived from our modeling results (Section II.B.3.c).
@@
The color coding in the shielding configuration regions represents magnetic field suppression efficiency in combination with system complexity: green denotes configurations with high field suppression and relatively simple designs, while red indicates systems that achieve high suppression but with excessive complexity. Additionally, the column colors specify the positioning of IR-absorbing layers within each shielding system. Together, these visual elements facilitate identification of optimal shielding configurations for SQCs. We would like to emphasize that the frequency of shielding systems mention in literature does not necessarily correlate with optimal performance. The subsequent analysis in this section aims to justify the prevalence of specific shielding configurations in the literature based on the presented in this paper simulation results.
IR shielding is typically implemented using a single absorbing layerPlace et al. (2021); Bal et al. (2024); Livingston et al. (2022) (Figure 14a), with only 10% of cases employing double-layer configurations Bengtsson (2020); Burnett et al. (2019) (Figure 14b). Our modeling results confirm the effectiveness of single-layer shielding positioned immediately adjacent to the sample holder (Figure 6), leading us to highlight these optimal configurations in sea-green in Figure 13b. Computational results further reveals the critical role of the sample holder’s interior surface in IR protection for SQCs. While literature examples of this implementation remain limited Kreikebaum et al. (2016) (Figure 14c), we have highlighted these cases through light-blue coding in Figure 13b.
Figure 14: Shielding systems with various IR shields: a) single-layer IR shielding combined with superconducting shieldPlace et al. (2021); b) a double-layer IR shielding: shields around the sample holder and on the lower cryostat stageBurnett et al. (2019); c) a double-layer IR shielding: absorbing coating on inner surface of the sample holder and IR shield around itKreikebaum et al. (2016).
Figure 15: Shielding systems with various superconducting shielding configurations: a) a single-layer superconducting shielding following after the sample holderYan et al. (2016); b) a single-layer superconducting shielding on copper base and following after IR absorbing layerGrünhaupt et al. (2018); c) a double-layer superconducting shielding: shields on the second and fifth place from the sample holderBengtsson (2020).
Figure 16: Shielding systems with various magnetic shielding configurations: a) a single-layer Amumetal shieldingReagor (2015); b) a double-layered Cryophy shieldingAsaad et al. (2016); c) a triple-layer $\mu$-metal shieldingJanzen et al. (2023).
@@
Approximately in 80% of considered works IR shielding is located around the sample holder Schwartz (2016); Gyenis et al. (2021); Jin et al. (2015); Serniak (2019) (Figure 14a). In other cases it surround the lowest stage of the fridge Burnett et al. (2019); Bengtsson (2020) as shown in Figure 14b . IR shields are typically positioned immediately after the sample holder Serniak (2019); Dickel et al. (2018). However, in multilayered shielding systems, they can be located at second Burnett et al. (2019) (Figure 14c) or even fourth Bengtsson (2020) shielding layer. Modeling results show that the multilayer IR shielding have the same effect as a single-layer one.
Double-layer configurations are implemented in approximately 20% of superconducting shielding systemsSingh et al. (2025); Bengtsson et al. (2020); Warren et al. (2023). This is well explained by the simulation results, according to which the number of superconducting shielding layers does not affect the shielding quality. These findings are further supported by the modeling results presented in previous sections. Our analysis demonstrates that a single-layer aluminum shield provides adequate protection against external magnetic field fluctuations for SQCs. As evident from Figure 13, approximately in a half of the examined shielding systems utilize superconducting sample holders, which simultaneously function as effective shielding components. This integrated approach represents a significant simplification of the overall shielding architecture.
@@
The location of the superconducting shields may vary, most of them directly surround the holder Jin et al. (2015), while the others located either at lower Burnett et al. (2019); Bengtsson (2020) or penultimate stage of the fridge Kreikebaum et al. (2016). Due to the requirement of the magnetic field to be uniformly distributed at the sample holder place, it is preferable to close the superconducting shield by the lid.
@@
Superconducting shields can be positioned either immediately adjacent to the sample holderYan et al. (2016); Oh et al. (2024); Eddins et al. (2019) (Figure 15a), or following the IR shielding layerGrünhaupt et al. (2018); Dickel et al. (2018); Jin et al. (2015) (Figure 15b). In more complex systems comprising four or more shields, superconducting layers may occupy the third or even fifth position relative to the sample holderBengtsson (2020) (Figure 15c). However, our modeling results showed that multilayered shielding system do not significantly increase shielding efficiency, so a more compact design may be used.
Our analysis of the literature reveals that single-layer magnetic shielding Reagor (2015); Braumüller et al. (2017); Dodge et al. (2023) (Figure 16a) predominates, being employed in 77% of surveyed studies. Two-layer configurations Asaad et al. (2016); Warren et al. (2023); Norris et al. (2024); Zwanenburg et al. (2025) (Figure 16b) were implemented in 16% of cases, while only a single research group reported utilizing a three-layer shielding system Janzen et al. (2023); Buks et al. (2020) (Figure 16c). As it was previously demonstrated, the combination of a superconducting shield inside and a ${\mu}$-metal shield outside provides the most effective protection against magnetic fields.
@@
As well as the other types of the shields, magnetic shields are typically placed around the holder Burnett et al. (2019); Kelly (2015); Sung (2018). At the same time, there are more options for the location of magnetic shields within the fridge, such as the lower stage Jin et al. (2015), 2.7 $K$ Kreikebaum et al. (2016) or around the fridge itself Moghaddam et al. (2019).
@@
Modeling results demonstrate that effective shielding requires fully enclosed configurations. Therefore, the shield must completely surround the sample holder to achieve optimal performance. In case the magnetic shield is the only one shield used in a system, it typically follows right after the sample holder Dodge et al. (2023); Chapman et al. (2023); Oh et al. (2024). In other cases, the magnetic shield may be placed outside of the other shields, such as IR shields Chou (2018); Wu et al. (2021) and superconducting shields Monroe et al. (2021a); Asaad et al. (2016). In multilayered magnetic shielding systems, ${\mu}$-metal shields can be incorporated between either superconducting or IR shielding layersBengtsson (2020); Burnett et al. (2019). As it was previously mentioned, for optimal shielding performance, the superconducting shield should be placed inside the ${\mu}$-metal one.
@@
Most considered shielding systems meet our requirements for the shielding. Requirements description for the shields based on the modeling results make it possible to identify excessiveness in the shielding system and make it lighter and more compact.
@@ III MICROWAVE LINES FILTERING
For high performance of SQC it is necessary to filter the frequencies out of band for all the types of lines listed above. Insufficient filtering of frequencies close to the qubit ones can lead to parasitic single-qubit rotations, degrading the single qubit gate fidelities, as well as undesirable population of possible coupled with qubits parasitic modes, which can lead to additional qubits decoherence. Insufficient filtering of IR frequencies (above 90 $GHz$) can lead to cooper-pairs breaking in superconductor thin film and also degrade the qubit coherence. In this article we focus only on the different filter types used for all lines listed above and made some recommendations for its filtering. The purpose of other microwave components, used in control lines, such as attenuators, circulators, isolators etc., is described in detail in Ref.Krinner et al. (2019).
@@ III.1 Requirements for filters
Filtering is usually used to protect SQCs from stray electromagnetic fields and IR radiation that can propagate in microwave lines. For this purpose, various types of microwave filters covering the SQC operating range (typically, 4-8 $GHz$) are usually used, namely Low-pass filter – LPFArute et al. (2019); Premkumar et al. (2021); Serniak (2019); Moghaddam et al. (2019); Bengtsson (2020), High-pass filter - HPFBurnett (2013); Kudra et al. (2020) or Band-pass filter - BPFBuks et al. (2020); Lisenfeld et al. (2019); Mundhada (2020). This filters are designed to cut off a portion of the signal spectrum outside the denoted passband, suppressing most of the stray electromagnetic fields. For suppressing IR radiation, a special IR filters are usually used. The working principle of such filters based on IR radiation power loss due to eddy currents excitation in filter filler material Wang et al. (2014).
It is necessary to use microwave filters with IR ones, as they cut-off different parts of the signal spectrum. The main requirement for IR filters is as low as possible attenuation in the SQC operating frequency range, smooth transmission characteristic and at the same time high enough attenuation (no less than 20 $dB$) at higher frequencies (above 20 $GHz$).
@@ III.2 IR filters
In works, IR filters may have different designations (IR filter Chou (2018); Opremcak et al. (2021); Kelly (2015); Schneider et al. (2019), lossy filter Schwartz (2016); Tan et al. (2015), thermalization filter Kreikebaum et al. (2016), etc.), but all of them are designed for the same purpose: to suppress stray IR radiation in signal lines. The filling material for such filters (see Figure 16) typically consists of an epoxy resin containing fine particles of carbonyl iron (Eccosorb CR-110 Place et al. (2021); Narla (2017); Reagor (2015), or, less frequently, CR-124 Krinner et al. (2019)), and metal powder Barends et al. (2011) (mainly copper Colless et al. (2018)). Other possible fillers include epoxy resin based on aluminum oxide (Stycast) mixed with bronze powder de Visser (2014), or without it Lisenfeld et al. (2019). Specific fillers such as carbon nanotubes may also be used Moghaddam et al. (2019).
There are no studies with a direct comparison of the IR filters efficiency with different types of filling material. However, in considered works, Eccosorb CR-110 epoxy is the most widely used in IR filters (see Figure 17).
@@
The attenuation coefficient of Eccosorb CR-110 exhibits a frequency-dependent increase from 5.4 $dB/cm$ at 100 $GHz$ to 90 $dB/cm$ at 1 $THz$Connolly et al. (2024). Due to its enhanced effectiveness at higher frequencies, even compact Eccosorb filters can effectively suppress Cooper-pair-breaking photons above 100 $GHz$Connolly et al. (2024).
The optimal filter length should be determined both by the required attenuation at the Cooper pair-breaking threshold frequency (e.g., $\sim$100 $GHz$ for 100 $nm$ thick aluminum films), and the maximum permissible attenuation at the qubit control frequency. For instance, in workConnolly et al. (2024) was shown that a 20 $mm$ long filter providing 0.5 $dB$ attenuation at 5 $GHz$ and ensure a charge parity switching rate of 10 $s^{-1}$, corresponding to a relaxation time limit of 100 $ms$.
@@
IR filters are employed in all types of signal lines of cryogenic setup. In the following sections we will consider importance of IR filter using in different line types.
III.3 Drive line filtering (XY control)